# Extended Circular Nim

### v1

@@ Abstract
Circular nim $CN(m,k)$ is a variant of nim, in which there are $m$ piles of tokens arranged in a circle and each player, in their turn, chooses at most $k$ consecutive piles in the circle and removes an arbitrary number of tokens from each pile. The player must remove at least one token in total. For some cases of $m$ and $k$, closed formulas to determine which player has a winning strategy have been found. Almost all cases are still open problems. In this paper, we consider a variant of circular nim, extended circular nim. In extended circular nim $ECN(m_{S},k)$, there are $m$ piles of tokes arranged in a circle. $S$ is a set of positive integers less than or equal to half of $m$. Each player, in their turn, chooses at most $k$ piles selected every $s$-th pile in a circle for an $s\in S$. We find some closed formulas to determine which player has a winning strategy for the cases where the number of piles is no more than eight.
@@ 1 Introduction
In this paper, we consider a generalization of a ruleset, circular nim. In circular nim $CN(m,k)$, there are $m$ piles of tokens arranged in a circle. Each player, in their turn, chooses at most $k$ consecutive piles in the circle and removes an arbitrary number of tokens from each pile. The player must remove at least one token in total. This game has been studied in some articles and winning positions for the previous player are characterized for some cases of $(m,k)$.
In this paper, we consider extended circular nim $ECN(m_{S},k)$, a generalization of circular nim. In $ECN(m_{S},k)$, there are $m$ piles of tokes arranged in a circle. $S$ is a set of positive integers less than or equal to half of $m$. Each player, in their turn, chooses at most $k$ piles selected every $s$-th pile in a circle for an $s\in S$. That is, $ECS(m_{\{1\}},k)$ is the same as $CS(m,k).$
@@ 1.1 Combinatorial game theory
In order to study games such as extended circular nim, we need to introduce some background of combinatorial game theory (CGT). Combinatorial game theory studies two-player perfect information games with no chance. In this paper, we consider impartial games, that is, in every position, both players have the same sets of options. We also assume that the games are short, that is, the game must end in a finite number of moves, and under normal play convention, that is, the player who cannot move is the loser.
With these assumptions, the player who has a winning strategy is uniquely determined for each position. A position in which the next player has a winning strategy is called an $\mathcal{N}$-position and a position in which the previous player has a winning strategy is called a $\mathcal{P}$-position.
It is well-known that if all positions of an impartial ruleset are separated into two sets $P$ and $N$, which satisfy the following properties, then $P$ and $N$ are the set of $\mathcal{P}$-positions and $\mathcal{N}$-positions, respectively:
@@
  • For any position in $P$, all the options of the position are in $N$.
  • For any position in $N$, at least one option of the position is in $P$.
We will use this fact for some proof.
@@ 1.2.1 Nim and Moore’s nim
A position in nim is denoted as $(n_{0},n_{1},\ldots,n_{m-1})$ when there are $m$ piles and each pile has $n_{0},n_{1},\ldots,n_{m-1}$ tokens. Let $\oplus$ be the exclusive OR operator for binary notation.
By using the following theorem, we can easily determine which player has a winning strategy in nim.
@@
When $k=1,$ Moore’s nim is the original nim. Thus, Moore’s nim is a generalization of nim. The $\mathcal{P}$-positions of this ruleset are characterized by the following theorem.
@@
A position $(n_{0},n_{1},\ldots,n_{m-1})$ in Moore’s nim ${\rm MN}(m,k)$, is a $\mathcal{P}$-position if and only if when each $n_{i}$ is expanded in the binary form, the summands of each digit contain a multiple of (k + 1) number of 1s.
This theorem is a generalization of Theorem 1.
Note that when $k=m-1$, a position $(n_{0},n_{1},\ldots,n_{m-1})$ in Moore’s nim ${\rm MN}(m,m-1)$ is a $\mathcal{P}$-position if and only if $n_{0}=n_{1}=\cdots=n_{m-1}$.
@@ 1.2.2 Simplicial nim
Let $V$ be a finite set of vertices. A simplicial complex $\Delta$ on $V$ is a subset of $2^{\mid V\mid}$ which satisfies the following:
  • For any element $v\in V,$$\{v\}\in\Delta$
@@
An element of $\Delta$ is termed a face of $\Delta$. Ehrenborg and Steingrímsson studied nim on simplicial complexes or simplicial nim [4]. Let $\Delta$ be a simplicial complex of a finite set $V$. Each vertex has several tokens. In nim on $\Delta$, the players, in turn, choose a non-empty face $F$ in $\Delta$ and arbitrarily remove any positive number of tokens from all vertices in $F$. The original nim and moore’s nim can be considered as special cases of this game. Ehrenborg and Steingrímsson did not characterize the $\mathcal{P}$-positions of nim on $\Delta$ without restrictions but discovered good constructions and characterized $\mathcal{P}$-positions for certain cases. For example, they proved that circuits, which are subsets of the vertex set and not faces of the simplicial complex, but all proper subsets of them are faces, can be used for characterizing $\mathcal{P}$-positions. Horrocks [6] and Penn [10] also researched this topic.
@@
Let $V=(v_{0},v_{1},v_{2})$ and $\Delta=\{\{v_{0}\},\{v_{1}\},\{v_{2}\},\{v_{0},v_{1}\}\}.$ Consider the position $A=(3,4,5)$ of nim on $\Delta$. In this case, the position $B=(1,2,5)$ is an option of $A$ because $\{v_{0},v_{1}\}$ is a face of $\Delta$ but the position $C=(3,3,3)$ is not an option of $A$ because $\{v_{1},v_{2}\}$ is not a face of $\Delta$.
@@ 1.2.3 Circular nim
Let $m$ and $k$ be positive integers with $k\leq m$. Circular nim ${\rm CN}(m,k)$ is nim on a simplicial complex $\Delta$, where $V=\{v_{0},\ldots,v_{m-1}\}$ and $\Delta=\bigcup_{i=0}^{m-1}\bigcup_{j=i}^{i+k-1}\{\{v_{i},v_{((i+1)\bmod m)},\ldots,v_{(j\bmod m)}\}\}$.
That is, in circula nim, there are $m$ piles of tokens arranged in a circle and each player, in their turn, chooses at most $k$ consecutive piles in the circle and removes an arbitrary number of tokens from each pile, where in total at least one token must be removed [^2] Note that in circular nim and extended circular nim defined in the next section, a pile is considered to be still existing when counting $k$ consecutive piles even if all tokens on the pile are removed. In contrast, in shrinking circular nim, a variant of circular nim discussed in [9], an empty pile is ignored.. When $k=1$, ${\rm CN}(m,k)$ is the same ruleset as $m$-pile nim. Also, when $k=m-1$, ${\rm CN}(m,k)$ is the same ruleset as Moore’s nim ${\rm MN}(m,m-1)$. Furthermore, when $k=m$, the player can select all of the piles; thus we can consider this ruleset as a single-pile nim. Therefore, these cases are trivial and other cases are studied in [4, 6, 2, 3] as follows. For any $M=(n_{0},\ldots,n_{m-1})$, we say $M\in_{\circlearrowleft}P$ if there exists $i(<m)$ such that $(n_{i},n_{((i+1)\bmod m)},\ldots,n_{((i+m-1)\bmod m)})\in P$ or $(n_{i},n_{((i+m-1)\bmod m)},\ldots,n_{((i+1)\bmod m)})\in P$.
@@
  • Consider a position $M=(n_{0},n_{1},n_{2},n_{3},n_{4})$ in ${\rm CN}(5,2)$. Then $M$ is a $\mathcal{P}$-position if and only if $M\in_{\circlearrowleft}P=\{(n_{0},n_{1},n_{2},n_{3},n_{4})\mid n_{0}=\max(M),n_{0}+n_{1}=n_{2}+n_{3},n_{1}=n_{4}\}$.
  • Consider a position $M=(n_{0},n_{1},n_{2},n_{3},n_{4})$ in ${\rm CN}(5,3)$. Then $M$ is a $\mathcal{P}$-position if and only if $M\in_{\circlearrowleft}P=\{(n_{0},n_{1},n_{2},n_{3},n_{4})\mid n_{0}=0,n_{1}=n_{2}+n_{3}=n_{4}\}.$
  • Consider a position $M=(n_{0},n_{1},n_{2},n_{3},n_{4},n_{5})$ in ${\rm CN}(6,3)$. Then $M$ is a $\mathcal{P}$-position if and only if $M\in_{\circlearrowleft}P=\{(n_{0},n_{1},n_{2},n_{3},n_{4},n_{5})\mid n_{0}+n_{1}=n_{3}+n_{4},n_{1}+n_{2}=n_{4}+n_{5}\}$.
  • Consider a position $M=(n_{0},n_{1},n_{2},n_{3},n_{4},n_{5})$ in ${\rm CN}(6,4)$. Then, $M$ is a $\mathcal{P}$-position if and only if $M\in_{\circlearrowleft}P=\{(n_{0},n_{1},n_{2},n_{3},n_{4},n_{5})\mid n_{0}=\min(M),n_{0}+n_{1}=n_{3}+n_{4},n_{1}+n_{2}=n_{4}+n_{5},n_{0}\oplus n_{2}\oplus n_{4}=~{}0\}$.
  • Consider a position $M=(n_{0},n_{1},n_{2},n_{3},n_{4},n_{5},n_{6})$ in ${\rm CN}(7,4)$. Then, $M$ is a $\mathcal{P}$-position if and only if $M\in_{\circlearrowleft}P=P_{1}\cup P_{2}\cup P_{3}\cup P_{4}$, where
@@
  • Consider a position $M=(n_{0},n_{1},n_{2},n_{3},n_{4},n_{5},n_{6},n_{7})$ in ${\rm CN}(8,6)$. Then, $M$ is a $\mathcal{P}$-position if and only if $M\in_{\circlearrowleft}P=\{(n_{0},n_{1},n_{2},n_{3},n_{4},n_{5},n_{6},n_{7})\mid n_{0}=0,n_{1}=n_{2}+n_{3}=n_{5}+n_{6}=n_{7},n_{4}=\min(\{n_{1},n_{2}+n_{6}\})\}$
For the other cases of circuar nim, it is still open problems to find a closed formula for determining the set of $\mathcal{P}$-positions.
@@ 1.3.1 Disjunctive sum
$\{G^{\prime}+H,G+H^{\prime}\mid G^{\prime}\text{ is an option of }G,H^{\prime}\text{ is an option of }H\}.$
@@
[…]

### v2

@@ Abstract
Circular nim ${\rm CN}(m,k)$ is a variant of nim, in which there are $m$ piles of tokens arranged in a circle and each player, in their turn, chooses at most $k$ consecutive piles in the circle and removes an arbitrary number of tokens from each pile. The player must remove at least one token in total. For some cases of $m$ and $k$, closed formulas to determine which player has a winning strategy have been found. Almost all cases are still open problems. In this paper, we consider a variant of circular nim, extended circular nim. In extended circular nim ${\rm ECN}(m_{S},k)$, there are $m$ piles of tokens arranged in a circle. $S$ is a set of positive integers less than or equal to half of $m$. In each turn, a player chooses an integer $s\in S$. Then the player selects at most $k$ piles among those located every $s$-th position on the circle, and removes an arbitrary number of tokens from each selected pile. We show some closed formulas to determine which player has a winning strategy for the cases where the number of piles is no more than eight, and for a few generalized cases.
@@ 1 Introduction
In this paper, we consider a generalization of a ruleset, circular nim. In circular nim ${\rm CN}(m,k)$, there are $m$ piles of tokens arranged in a circle. Each player, in their turn, chooses at most $k$ consecutive piles in the circle and removes an arbitrary nonnegative number of tokens from each pile. The player must remove at least one token in total. This game has been studied in some articles, and winning positions for the previous player are characterized for some cases of $(m,k)$. Figure 2 shows an example of a move in circular nim ${\rm CN}(6,3)$.
Figure 1: Move in circular nim ${\rm CN}(6,3)$.
Figure 2: Move in extended circular nim ${\rm ECN}(6_{\{1,3\}},2)$.
In this paper, we consider extended circular nim ${\rm ECN}(m_{S},k)$, a generalization of circular nim. In ${\rm ECN}(m_{S},k)$, there are $m$ piles of tokens arranged in a circle. $S$ is a set of positive integers less than or equal to half of $m$. Each player, in their turn, chooses at most $k$ piles selected every $s$-th pile in a circle for an $s\in S$. That is, ${\rm ECN}(m_{\{1\}},k)$ is the same as ${\rm CN}(m,k).$ Figure 2 shows an example of a move in extended circular nim ${\rm ECN}(6_{\{1,3\}},2)$.
@@ 1.1 Combinatorial game theory
In order to study games such as extended circular nim, we need to introduce some background of combinatorial game theory (CGT). Combinatorial game theory studies two-player perfect information games with no chance. In this paper, we use some theories of impartial rulesets, that is, in every position, both players have the same sets of options. We also assume that the games are short, that is, the game must end in a finite number of moves, and under normal play convention, that is, the player who cannot move is the loser.
With these assumptions, the player who has a winning strategy is uniquely determined for each position. A position in which the next player has a winning strategy is called an $\mathcal{N}$-position, and a position in which the previous player has a winning strategy is called a $\mathcal{P}$-position.
It is well-known that if all positions of an impartial ruleset are separated into two sets $P$ and $N$, which satisfy the following properties, then $P$ and $N$ are the sets of all $\mathcal{P}$-positions and all $\mathcal{N}$-positions, respectively:
@@
  • For every position in $P$, all options of the position are in $N$.
  • For every position in $N$, at least one option of the position is in $P$.
We will use this fact for some proofs. For the details of combinatorial game theory, see [15].
@@ 1.2.1 Nim and Moore’s nim
A position in nim is denoted as $(n_{0},n_{1},\ldots,n_{m-1})$ when there are $m$ piles and each pile has $n_{0},n_{1},\ldots,n_{m-1}$ tokens. Let $\oplus$ be the exclusive OR operator for binary notation. By using the following theorem, we can easily determine which player has a winning strategy in nim.
@@
When $k=1,$ Moore’s nim is the original nim. Thus, Moore’s nim is a generalization of nim. The $\mathcal{P}$-positions of this ruleset are characterized by the following theorem, which generalizes Theorem 1.
@@
A position $(n_{0},n_{1},\ldots,n_{m-1})$ in Moore’s nim ${\rm MN}(m,k)$ is a $\mathcal{P}$-position if and only if when each $n_{i}$ is expanded in the binary form, the summand of each digit contains a multiple of $(k+1)$ number of $1$s.
Note that when k=m1k=m-1, a position $(n_{0},n_{1},\ldots,n_{m-1})$ in Moore’s nim ${\rm MN}(m,m-1)$ is a $\mathcal{P}$-position if and only if $n_{0}=n_{1}=\cdots=n_{m-1}$.
@@ 1.2.2 Simplicial nim
Let $V$ be a finite set of vertices. A simplicial complex $\Delta$ on $V$ is a subset of $2^{V}$ which satisfies the following:
  • For every element $v\in V,$ $\{v\}\in\Delta$
@@
An element of $\Delta$ is termed a face of $\Delta$. Ehrenborg and Steingrímsson studied nim on simplicial complexes or simplicial nim in [8]. Let $\Delta$ be a simplicial complex on a finite set $V$. Each vertex has several tokens. In nim on $\Delta$, the players, in turn, choose a non-empty face $F$ in $\Delta$ and arbitrarily remove any positive number of tokens from all vertices in $F$. The original nim and moore’s nim can be considered as special cases of this game. Ehrenborg and Steingrímsson did not characterize the $\mathcal{P}$-positions of nim on $\Delta$ without restrictions but discovered good constructions and characterized $\mathcal{P}$-positions for certain cases. For example, they proved that circuits, which are subsets of the vertex set and not faces of the simplicial complex, but all proper subsets of them are faces, can be used for characterizing $\mathcal{P}$-positions. Horrocks [10] and Penn [14] also studied this ruleset. Hypergraph nim, studied in [2, 3, 4], is known as a generalization of simplicial nim.
@@
Let $V=(v_{0},v_{1},v_{2})$ and $\Delta=\{\{v_{0}\},\{v_{1}\},\{v_{2}\},\{v_{0},v_{1}\}\}.$ Consider the position $A=(3,4,5)$ of nim on $\Delta$. In this case, the position $B=(1,2,5)$ is an option of $A$ since $\{v_{0},v_{1}\}$ is a face of $\Delta$ but the position $C=(3,3,3)$ is not an option of $A$ since $\{v_{1},v_{2}\}$ is not a face of $\Delta$. Figure 3 shows these legal and illegal moves.
Figure 3: A legal move and an illegal move in simplicial nim.
For the later discussion, we refer to the following Definitions 3, 4 and Theorem 3 from [8].
Definition 3.
Given a complex $\Delta$ with vertex set $V$, let $\mathbb{N}^{V}$ be the set of vectors indexed by $V$ whose entries are non-negative integers. Let ${\bf e}(v)$ be the $v$-th unit vector, that is, the vector whose $v$-th entry is $1$ and all other entries are 0. For a subset $A$ of $V$, let
${\bf e}(A)=\sum_{v\in A}{\bf e}(v).$
Note that using this notation, a position $G=(n_{0},\ldots,n_{m-1})$ can be written as
$G=\sum_{0\leq i<|V|}n_{i}\cdot{\bf e}(v_{i}).$
Definition 4.
Let $\Delta$ be a simplicial complex and $\mathcal{B}$ a collection of subsets of $\Delta$ such that the following Conditions (A), (B), and (C) are satisfied. Here, $A=B\sqcup C$ means $A=B\cup C$ and $B\cap C=\emptyset.$
  • (A)
    The empty set belongs to $\mathcal{B}$.
  • (B)
    Suppose that $F$ is a face of $\Delta$, that $B,B^{\prime}\in\mathcal{B}$ and that $B^{\prime}=F\sqcup B$. Then $F$ is the empty face.
  • (C)
    Let $F$ be a face of $\Delta$ and let $S$ be a subset of $V$. Then there exist faces $K$ and $G$ of $\Delta$, with $K\subseteq F\subseteq G$, such that $G-F\subseteq S$ and $(S-G)\sqcup K\in\mathcal{B}$.
Then $\mathcal{B}$ is said to be a nim-basis for $\Delta$. A simplicial complex which has a nim-basis is said to be nim-regular.
Theorem 3.
Assume that $\Delta$ is a nim-regular simplicial complex with nim-basis $\mathcal{B}$. Then, a position $G$ is a $\mathcal{P}$-position on $\Delta$ if and only if
$G=\sum_{i\geq 0}2^{i}\cdot{\bf e}(A_{i}),$
where $A_{i}$ belongs to $\mathcal{B}$ for all $i\geq 0$.
@@ 1.2.3 Circular nim
Let $m$ and $k$ be positive integers with $k\leq m$. Circular nim ${\rm CN}(m,k)$ is nim on a simplicial complex $\Delta$, where $V=\{v_{0},\ldots,v_{m-1}\}$ and $\Delta$ is the set of all non-empty subset of $\bigcup_{i=0}^{m-1}\{v_{i},v_{((i+1)\bmod m)},\ldots,v_{((i+k-1)\bmod m)}\}$.
That is, in circula nim, there are $m$ piles of tokens arranged in a circle and each player, in their turn, chooses at most $k$ consecutive piles in the circle and removes an arbitrary number of tokens from each pile, where in total at least one token must be removed [^2] Note that in circular nim and extended circular nim defined in the next section, a pile is considered to be still existing when counting $k$ consecutive piles even if all tokens on the pile have been removed. In contrast, in shrinking circular nim, a variant of circular nim discussed in [13], an empty pile is ignored.. When $k=1$, ${\rm CN}(m,k)$ is the same ruleset as $m$-pile nim. Also, when k=m1k=m-1, ${\rm CN}(m,k)$ is the same ruleset as Moore’s nim ${\rm MN}(m,m-1)$. Furthermore, when $k=m$, the player can select all of the piles; thus, we can consider this ruleset as a single-pile nim. Therefore, these cases are trivial and other cases are studied in [8, 5, 10, 6, 7] as follows. For any $M=(n_{0},\ldots,n_{m-1})$, we say $M\in_{\circlearrowleft}P$ if there exists $i(<m)$ such that $(n_{i},n_{((i+1)\bmod m)},\ldots,n_{((i+m-1)\bmod m)})\in P$ or $(n_{i},n_{((i+m-1)\bmod m)},\ldots,n_{((i+1)\bmod m)})\in P$. That is, if a position in $P$ is equal to $M$, or it is the same position as $M$ after rotating or flipping it, then we say $M\in_{\circlearrowleft}P$.
@@
  • Consider a position $M=(n_{0},n_{1},n_{2},n_{3},n_{4})$ in ${\rm CN}(5,2)$. $M$ is a $\mathcal{P}$-position if and only if $M\in_{\circlearrowleft}P=\{(n_{0},n_{1},n_{2},n_{3},n_{4})\mid n_{0}=\max(M),n_{0}+n_{1}=n_{2}+n_{3},n_{1}=n_{4}\}$.
  • Consider a position $M=(n_{0},n_{1},n_{2},n_{3},n_{4})$ in ${\rm CN}(5,3)$. $M$ is a $\mathcal{P}$-position if and only if $M\in_{\circlearrowleft}P=\{(n_{0},n_{1},n_{2},n_{3},n_{4})\mid n_{0}=0,n_{1}=n_{2}+n_{3}=n_{4}\}.$
  • Consider a position $M=(n_{0},n_{1},n_{2},n_{3},n_{4},n_{5})$ in ${\rm CN}(6,3)$. $M$ is a $\mathcal{P}$-position if and only if $M\in_{\circlearrowleft}P=\{(n_{0},n_{1},n_{2},n_{3},n_{4},n_{5})\mid n_{0}+n_{1}=n_{3}+n_{4},n_{1}+n_{2}=n_{4}+n_{5}\}$.
  • Consider a position $M=(n_{0},n_{1},n_{2},n_{3},n_{4},n_{5})$ in ${\rm CN}(6,4)$. $M$ is a $\mathcal{P}$-position if and only if $M\in_{\circlearrowleft}P=\{(n_{0},n_{1},n_{2},n_{3},n_{4},n_{5})\mid n_{0}=\min(M),n_{0}+n_{1}=n_{3}+n_{4},n_{1}+n_{2}=n_{4}+n_{5},n_{0}\oplus n_{2}\oplus n_{4}=~0\}$.
  • Consider a position $M=(n_{0},n_{1},n_{2},n_{3},n_{4},n_{5},n_{6})$ in ${\rm CN}(7,4)$. $M$ is a $\mathcal{P}$-position if and only if $M\in_{\circlearrowleft}P=P_{1}\cup P_{2}\cup P_{3}\cup P_{4}$, where
@@
  • Consider a position $M=(n_{0},n_{1},n_{2},n_{3},n_{4},n_{5},n_{6},n_{7})$ in ${\rm CN}(8,6)$. $M$ is a $\mathcal{P}$-position if and only if $M\in_{\circlearrowleft}P=\{(n_{0},n_{1},n_{2},n_{3},n_{4},n_{5},n_{6},n_{7})\mid n_{0}=0,n_{1}=n_{2}+n_{3}=n_{5}+n_{6}=n_{7},n_{4}=\min(\{n_{1},n_{2}+n_{6}\})\}$.
For the other cases of circuar nim, it is still an open problem to find closed formulas for determining the set of $\mathcal{P}$-positions.
@@ 1.3.1 Disjunctive sum
[…]
@@
Let $G$ be a position of an impartial ruleset. The Sprague-Grundy value of $G$ is denoted by $\mathcal{G}(G)$ and
@@
For a position $(n_{0},n_{1},\ldots,n_{m-1})$ in nim it is easy to confirm that its Sprague-Grundy value is $n_{0}\oplus n_{1}\oplus\cdots\oplus n_{m-1}$. On the other hand, for Moore’s nim ${\rm MN}(m,k)$, no closed formula for calculating the Sprague-Grundy value in the general case is known.
However, for ${\rm MN}(m,m-1)$, a closed formula for calculating Sprague-Grundy values is shown as follows.
Theorem 5 ([11]).
Assume that $G=(n_{0},n_{1},\ldots,n_{m-1})$ is a position in ${\rm MN}(m,m-1)$ and $n_{0}=\min(\{n_{i}\mid 0\leq i\leq m-1\})$. We also let $t_{i}$ be $i$-th triangular number, that is, $t_{i}=\frac{i(i+1)}{2}$. Then, the Sprague-Grundy value of $G$ is
$\left\{\begin{array}[]{cc}t_{x}+((n_{0}-t_{x}-1)\bmod(x+1))&(\text{If }t_{x}<n_{0})\\ n_{0}+n_{1}+\cdots+n_{m-1}&(\text{Otherwise}),\end{array}\right$.
where $x=(n_{1}-n_{0})+\cdots+(n_{m-1}-n_{0})$.